OpenStudy (anonymous):

what is the gcf of 45 55 20

OpenStudy (anonymous):

5

OpenStudy (anonymous):

5 http://www.calculatorsoup.com/calculators/math/gcf.php

OpenStudy (jdoe0001):

\(\large \begin{cases} 45\to &3\cdot 3\cdot 5\\ 55\to &5\cdot 11\\ 20\to &2\cdot 2\cdot 5 \end{cases}\) what do you think?